OpenStudy (anonymous):

Please Help The solution set for -2x^2 + 12x = 0 is a. (0,12) b. (0,-12) c. (0,6) d. (0,-6)

OpenStudy (anonymous):

Can you find similar things to factor out in both terms?

OpenStudy (anonymous):

Try factoring out -2x

OpenStudy (anonymous):

\[-2x^2+12x=-2x(x-6)=0\]

OpenStudy (anonymous):

For these things to multiply together and equal to 0 one of them has to be =0, so either -2x=0 hence x=0, or x-6=0 hence x=6

OpenStudy (anonymous):

wait. I'm sorry its -2x^2+12x=0

OpenStudy (anonymous):

thats the same as you wrote